The electrolyte in automobile lead batteries is a 4.04 M sulfuric acid (H 2SO 4) solution. Calculate the mass percent of solute if the density of the solution is 1.161 g/mL. Express final answer in two decimal places.


1
Expert's answer
2021-09-09T05:50:02-0400

Solution


"1.161 \frac{g}{ml}\times{\frac{1000ml}{1L}}=1.161 \times 10^3"


4.04"mol\times {\frac{98.1gH_2SO_4}{1mol}}=396.32g"


Mass percent of the solute


"=\frac{396.32g}{1161g} \times100=34.14"


Mass percent of H2SO4 = 34.14%